# Scaling Latent Reasoning via Looped Language Models

### v1

@@ 1 Introduction
We explore a third pathway based on architectural innovation: achieving dynamic computation within a fixed parameter budget. This is accomplished by recursively reapplying shared parameters, where a group of weight-tied layers are iteratively reused during the forward pass. We call this the Looped Language Model (LoopLM). The LoopLM paradigm has gained wide attention and shown strong promise recently, originating with the seminal Universal Transformer [7]. Plenty of works on looped transformers [8, 9, 10, 11, 12, 13], recursive transformers [14], and latent reasoning frameworks [15, 16] have demonstrated the benefits of deeper computational processing on equivalent training data, indicating the potential benefits at a larger scale. These approaches share a common principle: through iterative reuse of weight-shared layers, computational scaling is decoupled into the depth of computation and the number of parameters involved.
The LoopLM design yields several advantages. First, LoopLM allows adaptive computation via a learned early exit mechanism: simple inputs may terminate after fewer recurrent steps, while complex ones naturally allocate more iterations, enabling flexible computational depth without increasing the parameter count. Moreover, unlike inference-time methods like CoT, LoopLM achieves scalability by deepening its internal computational graph rather than extending the output sequence, thereby avoiding excessive context growth. Finally, LoopLM potentially improves capacity per parameter and achieves stronger performance than standard transformers of comparable size when trained on the same data. While the prior studies have demonstrated the benefits of LoopLM at modest scales, there is little evidence that Looped Language Models can perform on par with non-looped architecture at practically meaningful scales. To this end, we ask the following question:
Does LoopLM exhibit more favorable scaling behavior (in loss, efficiency, safety, or capability growth),
compared to existing non-recursive transformer models?
In this work, we aim to answer this question affirmatively. We carefully investigate the LoopLM scaling trajectory to understand its point of saturation, illustrating that LoopLM represents a fundamentally more efficient path towards higher performance. While prior works have shown promise on modest scales (e.g., hundreds of billions of tokens), we study whether these benefits persist or even amplify when scaled to the multi-trillion token regimes of training that define SOTA foundation models. In particular, we study the mechanisms of LoopLM by further asking the following questions:
@@
    Does the recursive application of weight-shared layers enhance the LoopLM’s capabilities in a manner analogous to increasing the number of unshared layers?
@@
    Are the performance gains LoopLM monotonic in the number of loops? Are there more factors involved in the use of the adaptive computation of LoopLM that are different with prior empirical results in smaller scale experiments?
To resolve these uncertainties, we move beyond small-scale proofs of concept and undertake a comprehensive empirical investigation. Our results are presented below.
@@
In this work, we attempt to address the above research questions through a multi-faceted approach. We scale up the pretraining of LoopLM to a total of 7.7T tokens and thoroughly investigate the scaling behavior of LoopLM in various aspects. To achieve adaptive computation with LoopLM, we develop novel training objectives that enable computationally efficient recurrent computation while maintaining peak performance. We further conducted several controlled experiments on synthetic tasks to understand the mechanism of LoopLM superiority. Specifically, our contributions are:
@@
  • Entropy regularized adaptive computation. We develop an entropy regularized training objective with a uniform prior over exit steps that enables unbiased exploration of all recurrent depths. Moreover, we add a focused training stage for adaptive gates training, specifically to optimize the tradeoff between computation efficiency and loop performance, allowing learned adaptive allocation of recurrent steps based on input difficulty.
  • Mechanistic understanding via synthetic tasks. Using controlled experiments inspired by the physics of language models framework, we show that recurrence does not increase raw knowledge storage (approximately 2 bits per parameter for both looped and non-looped models) but dramatically enhances knowledge manipulation capabilities on tasks requiring fact composition and multi-hop reasoning.
  • Improved safety and faithfulness. LoopLM architectures reduce harmfulness on HEx-PHI [17], with safety improving as recurrent steps increase, including extrapolated steps. In contrast to CoT, our iterative latent updates yield a causally faithful reasoning process rather than post hoc rationalization.
Our study of LoopLM establishes the number of recursion as a third scaling axis beyond model size and data, and we publicly release the Ouro model family (1.4B and 2.6B parameters) to demonstrate the benefits of LoopLM at scale.
@@ 2 Related Works
We define notation and confine background only to what is needed for presenting our method. A standard $L$-layer Transformer is a composition of layer functions. It processes an input sequence of hidden states $H^{(0)}\in\mathbb{R}^{N\times d}$ through a series of unique layers, parameterized by $\theta_{1},\dots,\theta_{L}$:
$H^{(L)}=\text{TransformerLayer}_{\theta_{L}}\big(\dots\text{TransformerLayer}_{\theta_{2}}(\text{TransformerLayer}_{\theta_{1}}(H^{(0)}))\dots\big)$
In contrast, the Universal Transformer [7] replaces $L$ distinct layers with a recurrent application of a single Transformer block repeatedly for $T$ updates:
$H^{(t)}=\text{TransformerLayer}_{\theta}(H^{(t-1)}),\quad\text{for }t=1,\dots,T\text{, with }H^{(0)}\text{ as input}$
The core ideas of this architecture have resurfaced in recent literature, with recurrent-depth structures used to improve the efficiency and reasoning capabilities of modern LLMs. For example, Geiping et al. [15] adopts a “recurrent depth” to scale test-time computation in latent space. Similarly, Saunshi et al. [8] demonstrates that “looped transformers” can match the performance of much deeper non-looped models on reasoning tasks, formally connecting looping to the generation of latent thoughts. The approach is refined by converting standard models into “Relaxed Recursive Transformers” with a common base block while injecting unique LoRA adapters across recursive steps [14]. Similar concepts have emerged under different terms, such as “pondering” in continuous space [16] and “inner thinking” for adaptive computation [18]. More advanced variants, such as Mixture-of-Recursions [19] combine recursive parameter efficiency with adaptive, token-level routing.
Across all these works, from the original Universal Transformer to its modern descendants, this emerging line of architectures can be understood in two complementary ways. From one perspective, it behaves like a deep Transformer where the weights of all layers are tied. From another, iteration functions as latent reasoning, where the hidden states $H^{(1)},\dots,H^{(T)}$ form a latent chain of thought that progressively refines the representation to solve a task. Taken together, these results suggest that models can improve their ability to reason by reusing computation internally without having to increase parameter count, shifting scale to substance.
@@
Here, the LoopLM’s iteration is viewed as latent reasoning where each step is a non-verbal “thought” that refines the model’s internal representation. Empirically, increasing the number of recurrent steps improves performance on complex reaasoning tasks [15, 8]. Some models make this process explicit by feeding hidden states back into the input. Coconut inserts a “continuous thought” token, which is derived from the previous steps’s last-layer hidden state, so the model can “ponder” in a continuous latent space [24]. CoTFormer interleaves activations back into the input before reapplying this augmented sequence to the shared layers [25]. These explicit feedback loops contrast with implicit LoopLM variants, where the entire thought process is contained within the evolution of hidden states from $H^{(t-1)}$ to $H^{(t)}$. Thus, both Perspective 1 (model compression) and Perspective 2 (latent reasoning) leverage shared-parameter iteration to improve parameter efficiency, and are being explored for enhance reasoning and efficient sequence-length expansion (e.g., PHD-Transformer [26]).
@@ 3 Learning Adaptive Latent Reasoning with LoopLM
Figure 3: Overview of Looped Language Model (LoopLM) architecture. Left (Training): During training, the model applies a stack of NN shared-weight layers for $n$ recurrent steps ($R=1$ to $R=n$). At each recurrent step $i$, an exit gate predicts the probability $p_{i}$ of exiting, and a language modeling head $\mathcal{L}_{i}$ computes the task loss. The training objective combines the expected task loss across all steps with an entropy regularization term $H(p_{1},\ldots,p_{n})$ to encourage exploration of different computational depths. Right (Inference): At inference time, the model can exit early based on the cumulative distribution function (CDF) computed from exit probabilities. When $\text{CDF}_{i}=\sum_{k=1}^{i}p_{k}$ exceeds a threshold, the model terminates at step $i$, enabling adaptive computation that allocates more steps to harder inputs while maintaining efficiency on simpler ones. The dashed line indicates potential future steps that may be skipped through early exit.
In this section, we shall formally define the LoopLM architecture based on (causal) transformers and explain how we train LoopLMs that are able to perform latent reasoning with adaptive computation. Figure˜3 illustrates our architecture during both training and inference phases. Our objective is to let the model decide how many recurrent steps to use for each token and each example, spending less compute on easy inputs and more compute on hard inputs, without sacrificing accuracy when many steps are available.
@@ 3.1 LoopLM Architecture
Let $\mathrm{Tr}_{\theta}(\cdot):\mathbb{R}^{M\times d}\to\mathbb{R}^{M\times d}$ denote a causal transformer layer equipped with parameter $\theta$, with hidden dimension $d$ and input length $M$. Moreover we let $\mathrm{lmhead}(\cdot):\mathbb{R}^{d}\to\mathbb{R}^{|V|}$ denote the language-modeling head where $V$ is the vocabulary of tokens, and $\mathrm{emb}(\cdot):\mathbb{R}^{|V|}\to\mathbb{R}^{d}$ denote the embedding layer. A typical non-looped causal language model is defined by stacking $L$ layers as follows:
[…]
Let $t\in\mathbb{N},t\leq T_{\max}$ be the number of loop steps, which we also call the number of recurrent steps or recurrent depth, we define the looped language model $F^{(t)}$ by:
$F^{(t)}(\cdot)=\mathrm{lmhead}\circ\underbrace{H^{L}\circ H^{L}\circ\cdots\circ H^{L}}_{t\text{ times}}\circ\ \mathrm{emb}(\cdot).$ (1)
Apparantly $F^{(1)}\equiv F$ is the non-loop model. Given a sequence of tokens $x_{1:M}=(x_{1},x_{2},\dots,x_{M})$ of length $M$, a LoopLM model $F^{(t)}$ acts as a sequence to sequence model and produces $M$ tokens $y_{1:M}=(y_{1},y_{2},\dots y_{M})$ that is consistent with the causal dependency condition that $y_{1:m}=F^{(t)}(x_{1:m})$ and for any $m\leq M$. The next-token-prediction objective is simply the cross-entropy loss on all $y_{i},1\leq i\leq M$,
$\mathcal{L}^{(t)}=\mathbb{E}_{x_{1:M}}\bigg[\sum_{1\leq i\leq M-1}-\log\mathbf{Pr}(F^{(t)}(x_{1:i})=x_{2:i+1})\bigg],\quad\text{for }t\in[1,T_{\max}]$ (2)
where $\mathbf{Pr}(F^{(t)}(x_{1:i})=x_{i+1})$ is computed by taking a softmax over the $\mathrm{lmhead}$ output. The loss $\mathcal{L}^{(t)}$ measures the precision of a $t$-step LoopLM. Prior literature such as [12, 8] have shown that scaling up the loop count $t$ is beneficial for reasoning tasks in smaller scale. However, increasing the loop count $t$ costs computation, and not all language tokens are reasoning-heavy and need those computation to be predicted correctly. In fact, many tokens in the next-token-prediction tasks are simple or can be learned with high certainty [27, 28]. Thus it is crucial to spend the computation budget on the right tokens for pareto-optimal allocation between performance and efficiency. This is achieved by the gating mechanism we shall describe below.
@@ 3.2 Adaptive Computation via Gating Mechanism
Algorithm 1 Q-exit Early Stopping Criterion
1:Input $x$; maximum steps $T_{\max}$; threshold $q\in[0,1]$;
[…]
3:for $t=1$ to $T_{\max}$ do
4:  $\lambda_{t}\leftarrow\sigma(\mathrm{Gate}(F^{(t)}(x)))$
5:  $q_{\phi}(t|x)\leftarrow\lambda_{t}\cdot surv$
6:  $cdf\leftarrow cdf+p_{t}$ $\triangleright$ $\mathrm{CDF}(t\mid x)$
7:  if $cdf\geq q$ then
8:   return $t$ $\triangleright$ $t_{\mathrm{exit}}(x)$
9:  $surv\leftarrow surv\cdot(1-\lambda_{t})$
10:return $T_{\max}$ $\triangleright$ fallback if the threshold is never reached
[…]
$\text{CDF}(t|x)=\sum_{i=1}^{t}q_{\phi}(i|x)=\sum_{i=1}^{t}\lambda_{i}(x)\prod_{j=1}^{i-1}(1-\lambda_{j}(x))$
and $q_{\phi}(i|x)$ represents the probability of exiting exactly at step $i$ given input $x$. The product term $\prod_{j=1}^{i-1}(1-\lambda_{j})$ ensures that we can only stop at step $i$ if we did not exit at any earlier step $j<i$. For input $x$, we exit early when the CDF exceeds a set threshold $q\in[0,1]$:
$t_{\text{exit}}(x)=\min\{t:\text{CDF}(t|x)\geq q\}$
The threshold $q$ controls the compute-accuracy tradeoff, where lower values encourage earlier exits and higher values allow deeper computation. The gating function $\lambda_{t}$ shall be learned in the two-stage training:
  • Stage I: during pre-training, the gates $\lambda_{t}(\cdot)$ are learned by optimizing an entropy-regularized objective. In this stage, the exit distribution aims to widen the coverage of the optimal exit step.
  • Stage II: we focus on training only the gates in this stage. The objective exploits the pattern learned in stage I and sharpens the distribution by choosing the optimal exit step with a soft cross-entropy loss.
We shall immediately explain the two stages below.
@@ 3.3 Stage I: Learning An Entropy-Regularized Objective
To obtain a LoopLM that can exit early while maintaining performance, we need an objective that jointly optimizes accuracy and computational efficiency. Let $\mathcal{L}^{(t)}$ follow (2), then our training objective combines the next-token prediction loss with an entropy regularization term over $q_{\phi}$:
@@
$\mathcal{L}=\underbrace{\sum_{t=1}^{T_{\max}}q_{\phi}(t|x)\cdot\mathcal{L}^{(t)}}_{\text{expected task loss}}-\underbrace{\beta\cdot H(q_{\phi}(\cdot|x))}_{\text{entropy regularization}}\qquad H(q_{\phi}(\cdot|x))=-\sum_{t=1}^{T_{\max}}q_{\phi}(t|x)\log q_{\phi}(t|x)$ (3)
where $\mathcal{L}^{(t)}$ is the next token prediction loss at loop step $t$, and $H(q_{\phi}(\cdot|x))$ is the entropy of the exit step distribution. The hyperparameter $\beta$ controls the exploration-exploitation trade-off: larger $\beta$ encourages more uniform distributions (higher entropy), allowing the model to explore different depths, while smaller $\beta$ allows more concentrated distributions when the model is confident about the optimal depth for a given input.
Alternative perspective: variational inference with uniform prior.
@@
The entropy-regularized objective can be equivalently viewed through the lens of variational inference. If we treat the exit step as a latent variable $z$ with a prior distribution $\pi$, we can derive an Evidence Lower Bound (ELBO) objective:
$\mathcal{L}_{\text{ELBO}}=\sum_{t=1}^{T_{\max}}q_{\phi}(t\mid x)\cdot\mathcal{L}^{(t)}+\beta\cdot\mathrm{KL}\big(q_{\phi}(\cdot\mid x)\,\|\,\pi\big)$
When we choose a uniform prior $\pi_{t}=1/T_{\max},\forall t$, the KL divergence simplifies to:
$\mathrm{KL}(q_{\phi}(\cdot\mid x)\|\pi)=-H(q_{\phi}(\cdot\mid x))+\log T_{\max}$
@@
Since $\log T_{\max}$ is constant, minimizing the ELBO with a uniform prior is equivalent to our entropy-regularized objective. This connection reveals that our approach aligns with adaptive computation methods like PonderNet [30], which also optimize an ELBO objective for dynamic halting.
Why uniform prior?
While our formulation is similar to PonderNet [30], a critical difference lies in the choice of prior. PonderNet and other adaptive computation methods typically employ geometric priors:
$\pi_{t}^{\mathrm{geo}}=\frac{\lambda\,(1-\lambda)^{\,t-1}}{1-(1-\lambda)^{\,T_{\max}}},\quad t=1,\dots,T_{\max},\quad\lambda\in(0,1)$
Similarly, methods like Recurrent Depth [15] use Poisson-lognormal priors that also favor shallow computation. We argue that such priors conflate two distinct goals: (1) learning when to exit based on input difficulty, and (2) minimizing average computation cost. In fact, these priors contain a strong inductive bias toward shallow computation by placing more mass on early steps, explicitly encouraging the model to exit early. Therefore, they risk under-exploring deeper steps and may fail to fully exploit the benefits of recurrent depth.
In contrast, the uniform prior makes no assumptions about the optimal exit step distribution. It optimizes all depths uniformly, enabling the model to infer the computational requirements of different inputs without relying on predefined inductive biases. This is particularly important for complex reasoning tasks, where the optimal depth should be learned from the data rather than being constrained by the prior. We provide a detailed empirical validation of this choice against geometric priors in Appendix˜A.

### v2

@@ 1 Introduction
We explore a third pathway based on architectural innovation: achieving dynamic computation within a fixed parameter budget. This is accomplished by recursively applying shared parameters, where a group of weight-tied layers are iteratively reused during the forward pass. We call this the Looped Language Model (LoopLM). The design yields several advantages. First, LoopLM enables adaptive computation via a learned early-exit mechanism: simple inputs can terminate after fewer recurrent steps, while complex ones allocate more iterations. This decouples the compute depth from parameter count. Second, unlike inference-time methods such as CoT, LoopLM scales by deepening its internal computational graph rather than extending the output sequence, avoiding context-length bloat. Finally, LoopLM can improve capacity per parameter and outperform standard transformers of larger sizes when trained on the same data.
An extensive range of prior studies have explored LoopLM at modest scales [7, 8, 9, 10, 11, 12, 13, 14], from the seminal Universal Transformer [15] to recursive Transformers [16] and latent reasoning approaches [17, 18, 19]. Yet whether Looped Language Models translate into frontier-level gains at practically meaningful scales is unproven. To this end, we ask:
Does LoopLM exhibit more favorable scaling behavior (loss, efficiency, safety, capability growth),
compared to non-recursive transformer models?
We show the answer is yes. We characterize LoopLM’s scaling trajectory and saturation behavior, demonstrating that LoopLM offers a more efficient path to higher performance. These claims are evaluated under multi-trillion-token training regimes typical of SoTA foundation models, extending well beyond prior work. Beyond the empirical gains, we analyze the mechanisms behind these improvements by asking the following questions:
@@
    Does the recursive reuse of weights yield the capability gains typically obtained by increasing depth with unshared weights?
@@
    Are LoopLM’s gains monotonic in the number of loops? What are the factors that influence this?
@@
We address the above questions with a multi-faceted study. We scale LoopLM pre-training to 7.7T tokens and thoroughly investigate its scaling behavior across multiple axes. To enable adaptive computation, we introduce training objectives that enable computationally efficient recurrence while preserving peak performance. We also run controlled ablations to isolate the sources of LoopLM’s gains. Specifically, our contributions are:
@@
  • Entropy-regularized adaptive computation. Adaptive exits tend to collapse to shallow depths or overuse long loops. We avoid this with entropy-reguarization under a uniform prior over exit steps for unbiased depth exploration, followed by a focused training stage that tunes the compute-performance trade-off and allocates steps based on input difficulty.
  • Mechanistic understanding of recurrence. Using controlled experiments inspired by the physics-of-LMs framework, we find recurrence does not increase raw knowledge storage (approximately 2 bits per parameter for looped and non-looped models) but dramatically enhances knowledge manipulation capabilities on tasks requiring fact composition and multi-hop reasoning.
  • Improved safety and faithfulness. LoopLM reduces harmfulness on HEx-PHI [20], with safety improving as recurrent steps increase (including extrapolated steps). Compared to CoT, our iterative latent updates produce reasoning traces that are better aligned with final outputs, indicating greater causal faithfulness rather than post-hoc rationalization.
Our study establishes loop depth as a third scaling axis beyond model size and data, and we publicly release the Ouro model family (1.4B and 2.6B parameters) to demonstrate the benefits of LoopLM at scale.
@@ 2 Related Works
The core ideas of this architecture have resurfaced in recent literature, with recurrent-depth structures used to improve the efficiency and reasoning capabilities of modern LLMs. For example, Geiping et al. [17] adopts a “recurrent depth” to scale test-time computation in latent space. Similarly, Saunshi et al. [7] demonstrates that “looped transformers” can match the performance of much deeper non-looped models on reasoning tasks, formally connecting looping to the generation of latent thoughts. The approach is refined by converting standard models into “Relaxed Recursive Transformers” with a common base block while injecting unique LoRA adapters across recursive steps [16]. Similar concepts have emerged under different terms, such as “pondering” in continuous space [zeng2025pre-training] and “inner thinking” for adaptive computation [21]. More advanced variants, such as Mixture-of-Recursions [22] combine recursive parameter efficiency with adaptive, token-level routing.
Across all these works, from the original Universal Transformer to its modern descendants, this emerging line of architectures can be understood in two complementary ways. From one perspective, it behaves like a deep Transformer where the weights of all layers are tied. From another, iteration functions as latent reasoning, where the hidden states form a latent chain of thought that progressively refines the representation to solve a task. Taken together, these results suggest that models can improve their ability to reason by reusing computation internally without having to increase parameter count, shifting scale to substance.
@@
Here, the LoopLM’s iteration is viewed as latent reasoning where each step is a non-verbal “thought” that refines the model’s internal representation. Empirically, increasing the number of recurrent steps improves performance on complex reaasoning tasks [17, 7]. Some models make this process explicit by feeding hidden states back into the input. Coconut inserts a “continuous thought” token, which is derived from the previous steps’s last-layer hidden state, so the model can “ponder” in a continuous latent space [27]. CoTFormer interleaves activations back into the input before reapplying this augmented sequence to the shared layers [28]. These explicit feedback loops contrast with implicit LoopLM variants, where the entire thought process is contained within the evolution of hidden states from previous recurrent step to current. Thus, both Perspective 1 (model compression) and Perspective 2 (latent reasoning) leverage shared-parameter iteration to improve parameter efficiency, and are being explored for enhance reasoning and efficient sequence-length expansion (e.g., PHD-Transformer [29]).
@@ 3 Learning Adaptive Latent Reasoning with LoopLM
Figure 3: Overview of Looped Language Model (LoopLM) architecture. Left (Training): During training, the model applies a stack of $N$ layers repeatedly for $T_{max}$ recurrent steps. At each recurrent step $\ell$, an exit gate predicts the probability $p_{\ell}$ of exiting, and a language modeling head $\mathcal{L}_{\ell}$ computes the lanugage modeling loss. Right (Inference): At inference time, the model can exit early based on the accumulated exit probability.
In this section, we formally define the LoopLM architecture on causal Transformers and present our training scheme for adaptive latent reasoning. Figure˜3 illustrates the architecture in training and inference. Our goal is to let the model choose the number of recurrent steps per token and per example, spending less compute on easy inputs and more on hard inputs, without sacrificing accuracy when many steps are available.
@@ 3.1 LoopLM Architecture
Let $\mathrm{emb}(\cdot):\mathbb{R}^{|V|}\to\mathbb{R}^{d}$ be the token embedding; $\mathcal{T}_{\theta}(\cdot):\mathbb{R}^{M\times d}\to\mathbb{R}^{M\times d}$ a causal transformer layer parameterized by $\theta$ with hidden size $d$ and input length $M$, and $\mathrm{lmhead}(\cdot):\mathbb{R}^{d}\to\mathbb{R}^{|V|}$ be the unembedding layer with vocabulary size $V$. A non-looped LM stacks $L$ layers, where $\circ$ denotes function composition:
$F(\cdot):=\mathrm{lmhead}\circ\mathcal{M}^{L}\circ\mathrm{emb(\cdot)},\quad\qquad\mathcal{M}^{L}(\cdot):=\mathcal{T}_{\theta_{L}}\circ\cdots\circ\mathcal{T}_{\theta_{1}}(\cdot)$
Let $t\in\{1,\dots,T_{\max}\}$ be the number of loop steps (the number of recurrent steps or recurrent depth). The looped model $F^{(t)}$ reuses the same depth-$L$ layer stack $t$ times:
$F^{(t)}(\cdot)=\mathrm{lmhead}\circ\underbrace{\mathcal{M}^{L}\circ\mathcal{M}^{L}\circ\cdots\circ\mathcal{M}^{L}}_{t\text{ iterations}}\circ\ \mathrm{emb}(\cdot).$ (1)
Thus, $t=1$ yields the non-looped model $F^{(1)}\equiv F$. As shown in Figure 3 (Left), at each recurrent step $t$, the model produces a language modeling head output. We define the standard cross-entropy loss at single step $t$ as the loss, $\mathcal{L}^{(t)}$:
$\mathcal{L}^{(t)}=\mathbb{E}_{x_{1:M}}\Bigg[\sum_{\ell=1}^{M-1}-\log\,p^{(t)}_{\theta}\!\big(x_{\ell+1}\mid x_{1:\ell}\big)\Bigg],$ (2)
where $p^{(t)}_{\theta}(\cdot\mid x_{1:\ell})=\mathrm{softmax}\!\big(\mathrm{lmhead}(h^{(t)}_{\ell})\big)$, $x_{1:\ell}$ denotes the length-$\ell$ prefix of the input (tokens $1$ through ellell), and $h^{(t)}_{\ell}$ is the hidden state after $t$ loops at position $\ell$. Note that this is the individual loss for a single recurrent step. The total training objective, which combines all steps, is defined in the following sections.
Prior literature [11, 7] has shown that scaling up $t$ is beneficial for reasoning tasks. However, this increases computation, and not all tokens require many steps [30, 31]. Thus, it is crucial to spend the computation budget on the right tokens. This is achieved by the gating mechanism described in the next section.
@@ 3.2 Adaptive Computation via Gating Mechanism
To enable adaptive computation, we add an exit gate that runs in parallel with the LM head at each step $t\leq T_{\max}$ (Figure˜3). At each loop $t$, the gate outputs an instantaneous (per-step) exit probability
$\lambda_{t}(x)=\sigma\left(\mathrm{Linear}_{\phi}\left(h^{(t)}\right)\right)\in(0,1)$
where $h^{(t)}$ is the final-layer hidden state at step $t$ and $\phi$ are the gate parameters. We define
$S_{t}(x)=\prod_{j=1}^{t}\bigl(1-\lambda_{j}(x)\bigr),\qquad S_{0}(x)\equiv 1,$
as the survival, or the probability of not exiting in the first $t$ steps. The unnormalized probability of exiting first at step $t$ is then
$\tilde{p}_{t}(x)=\lambda_{t}(x)\,S_{t-1}(x),\qquad t=1,\dots,T_{\max}-1.$
To obtain a valid discrete distribution over exit steps, we assign the remaining mass to the final step:
$p_{\phi}(t\mid x)=\begin{cases}\tilde{p}_{t}(x),&t=1,\dots,T_{\max}-1,\\[3.0pt] S_{T_{\max}-1}(x),&t=T_{\max},\end{cases}\qquad\text{so that}\quad\sum_{t=1}^{T_{\max}}p_{\phi}(t\mid x)=1.$ (3)
Inference with Early Exit.
As illustrated in Figure 3 (Right), we infer an exit step from the learned exit distribution $\{p_{\phi}(t\mid x)\}_{t=1}^{T_{\max}}$, enabling efficient inference. The cumulative exit probability up to step $n$ is:
$\mathrm{CDF}(n\mid x)=\sum_{t=1}^{n}p_{\phi}(t\mid x)=1-\prod_{j=1}^{n}\bigl(1-\lambda_{j}(x)\bigr),\quad n<T_{\max},\qquad\mathrm{CDF}(T_{\max}\mid x)=1.$
Given a threshold $q\in[0,1]$, we terminate at the first step where the cumulative probability crosses $q$:
$t_{\mathrm{exit}}(x)=\min\{\,m\in\{1,\dots,T_{\max}\}\,\;:\;\mathrm{CDF}(m\mid x)\geq q\,\}.$
The threshold $q$ controls the compute–accuracy tradeoff: smaller $q$ favors earlier exits (less compute), while larger $q$ allows deeper computation. In practice, $q$ may be chosen globally, calibrated per task, or scheduled with a floor/ceiling on steps. This deterministic, quantile-based policy avoids sampling while remaining consistent with the learned distribution.
The gating parameters $\phi$ (and thus $p_{\phi}$ via $\{\lambda_{t}\}$) are learned in two stages:
  • Stage I: During pre-training, the gates are learned jointly with the LM by optimizing an entropy-regularized objective (Section˜3.3).
  • Stage II: We freeze the LM and fine-tune $\phi$ to sharpen $p_{\phi}$ (i.e., adjust depth allocation) without changing token-level predictions.
The complete training objective is described in the next section.
@@ 3.3 Stage I: Learning an Entropy-Regularized Objective
Under naive gradient descent on the next-token prediction loss, deeper loops typically reduce the single-step loss $\mathcal{L}^{(t)}$ from Equation˜2 up to some depth; beyond that, gains diminish and the gradients shift probability mass toward later steps. As $p_{\phi}$ concentrates on late steps, those steps receive more training signal and their losses drop further, which in turn pulls even more mass to the end. This self-reinforcement collapses $p_{\phi}$ onto $t=T_{\rm max}$. An entropy term penalizes collapse to the deepest step, maintaining enough spread in $p_{\phi}$ to reflect input difficulty.
Given the single-step loss $\mathcal{L}^{(t)}$ and the exit-step distribution $p_{\phi}(t\mid x)$ from Equation˜3, our training objective combines next-token prediction with entropy regularization:
@@
$\mathcal{L}=\underbrace{\sum_{t=1}^{T_{\max}}p_{\phi}(t\mid x)\,\mathcal{L}^{(t)}}_{\text{expected task loss}}-\underbrace{\beta\,H\!\left(p_{\phi}(\cdot\mid x)\right)}_{\text{entropy regularization}},\qquad H\!\left(p_{\phi}(\cdot\mid x)\right)=-\sum_{t=1}^{T_{\max}}p_{\phi}(t\mid x)\log p_{\phi}(t\mid x).$ (4)
Intuitively, the expected task loss weights each $\mathcal{L}^{(t)}$ by the probability of exiting at step $t$. The coefficient $\beta$ controls the exploration–exploitation trade-off: larger $\beta$ encourages higher-entropy (more exploratory) $p_{\phi}$, while smaller $\beta$ lets $p_{\phi}(t\mid x)$ place most of its mass on a specific step when the model is confident about the optimal depth.
Alternative perspective: variational inference with a uniform prior.
@@
The objective in Equation˜4 can be viewed as an Evidence Lower Bound (ELBO) loss where the exit step $z\in\{1,\dots,T_{\max}\}$ is a latent variable whose variational posterior is the learned exit distribution $p_{\phi}(z{=}t\mid x)$ and whose prior is $\pi(t)$. The negative ELBO is:
$\mathcal{L}_{\text{ELBO}}=\sum_{t=1}^{T_{\max}}p_{\phi}(t\mid x)\,\mathcal{L}^{(t)}\;+\;\beta\,\mathrm{KL}\!\big(p_{\phi}(\cdot\mid x)\,\|\,\pi(\cdot)\big).$
With a uniform prior $\pi_{t}=1/T_{\max}$, the KL becomes
$\mathrm{KL}\!\big(p_{\phi}(\cdot\mid x)\,\|\,\pi\big)=-H\!\left(p_{\phi}(\cdot\mid x)\right)+\log T_{\max},$
@@
so minimizing the ELBO is equivalent (up to the constant $\log T_{\max}$) to the objective in Equation˜4. This identifies the entropy term as a KL regularizer and clarifies that the expected loss marginalizes over exit steps, while also linking to adaptive-computation methods such as PonderNet [32], which also optimize an ELBO for dynamic halting.
Why a uniform prior?
Different priors encode different depth preferences. A geometric prior, as in Ref. [32], or Poisson-lognormal priors softly favor earlier halting [17], while a uniform prior is depth-unbiased. We adopt the uniform prior to decouple exit decisions driven by input difficulty from any global compute preference; the entropy term then prevents collapse to always using $T_{\rm max}$. Empirical comparisons with geometric priors are provided in Appendix Appendix˜A.